# 6DGS: Enhanced Direction-Aware Gaussian Splatting for Volumetric Rendering

### v1

@@ Abstract
Novel view synthesis has advanced significantly with the development of neural radiance fields (NeRF) and 3D Gaussian splatting (3DGS). However, achieving high quality without compromising real-time rendering remains challenging, particularly for physically-based ray tracing with view-dependent effects. Recently, N-dimensional Gaussians (N-DG) introduced a 6D spatial-angular representation to better incorporate view-dependent effects, but the Gaussian representation and control scheme are sub-optimal. In this paper, we revisit 6D Gaussians and introduce 6D Gaussian Splatting (6DGS), which enhances color and opacity representations and leverages the additional directional information in the 6D space for optimized Gaussian control. Our approach is fully compatible with the 3DGS framework and significantly improves real-time radiance field rendering by better modeling view-dependent effects and fine details. Experiments demonstrate that 6DGS significantly outperforms 3DGS and N-DG, achieving up to a 15.73 dB improvement in PSNR with a reduction of 66.5% Gaussian points compared to 3DGS.
Figure 1: Visualizations of volumetric rendering. Top-left: our 6DGS rendering results; bottom-right: physically-based rendering using ray tracing; right: comparison with 3DGS over the red crop regions.
@@ 1 Introduction
We validate our method through extensive experiments on two datasets: a custom dataset with physically based rendering using ray tracing, named the 6DGS-PBRT dataset, and a public dataset without strong view-dependent effects, i.e., the Synthetic NeRF dataset (Mildenhall et al., 2020). The 6DGS-PBRT dataset contains scenes with complex geometries, materials, and lighting conditions that exhibit strong view-dependent effects, making it suitable for testing our method’s capabilities in handling complex light interactions. Our results demonstrate that 6DGS significantly outperforms existing methods in both rendering quality and efficiency on the 6DGS-PBRT dataset and achieves comparable performance on the Synthetic NeRF dataset. Specifically, we achieve up to a 15.73 dB improvement in peak signal-to-noise ratio (PSNR) while using only 33.5% of the Gaussian points compared to 3DGS. These improvements underscore the effectiveness of our approach in capturing fine details and complex view-dependent effects in real-time rendering scenarios.
@@
  • Validation: We validate our 6DGS method on a custom physically-based ray tracing dataset (6DGS-PBRT), demonstrating its superiority in both image quality and rendering speed compared to existing approaches. Additionally, our evaluation on a public dataset showcases its generalizability to scenes without strong view-dependent effects.
@@ 2 Related Work
3D Gaussian splatting (3DGS) (Kerbl et al., 2023) has emerged as a significant advancement in computer graphics and 3D vision, achieving high-fidelity rendering quality while maintaining real-time performance. Numerous works have been proposed to improve rendering quality (Yu et al., 2024; Lu et al., 2024), rendering efficiency (Lee et al., 2024; Bagdasarian et al., 2024), and training optimization (Kheradmand et al., 2024; Höllein et al., 2024), as well as to explore applications (Kocabas et al., 2024; Zhou et al., 2024b; Niedermayr et al., 2024) and extensions (Charatan et al., 2024; Luiten et al., 2024; Wu et al., 2024; Tang et al., 2024) of 3DGS. For example, Mip-Splatting (Yu et al., 2024) introduces a Gaussian low-pass filter based on Nyquist’s theorem to address aliasing and dilation artifacts by matching the maximal sampling rate across all observed samples. Compact3D (Lee et al., 2024) applies vector quantization to compress different attributes into corresponding codebooks, storing the index of each Gaussian to reduce storage overhead. 3DGS-MCMC (Kheradmand et al., 2024) proposes densification and pruning strategies in 3DGS as deterministic state transitions of Markov Chain Monte Carlo (MCMC) samples instead of using heuristics. Many of these advancements and applications of 3DGS can potentially be applied to our 6DGS, which extends 3DGS by incorporating an additional directional component.
N-dimensional Gaussians.
@@
To enhance the 3D Gaussian representation, researchers have introduced other Gaussian representations for rendering. 2DGS (Huang et al., 2024) introduces a perspective-accurate 2D splatting process utilizing ray-splat intersection and rasterization to enhance geometry reconstruction. 3D-HGS (Li et al., 2024) proposes a 3D half-Gaussian kernel to improve performance without compromising rendering speed. 4DGS (Yang et al., 2024) proposes to approximate the underlying spatio-temporal 4D volume of a dynamic scene by optimizing a collection of 4D primitives. N-DG (Diolatzis et al., 2024) introduces N-dimensional Gaussians (N-DG) along with a high-dimensional culling scheme inspired by locality-sensitive hashing. Specifically, N-DG introduces a 10-dimensional Gaussian (10-DG) that incorporates geometry and material information such as world position, view direction, albedo, and roughness, as well as a 6-dimensional Gaussian (6-DG) that includes world position and view direction. Our 6DGS combines the strengths of 3DGS and N-DG for better representation and better adaptive control of Gaussians.
Gaussian-based Ray Tracing.
@@
Unlike most 3DGS methods that render Gaussians via rasterization, some approaches have proposed Gaussian-based ray tracing, generally at the cost of slower rendering speeds or even lower quality. For instance, Condor et al. (2024) models scattering and emissive media using mixtures of simple kernel-based volumetric primitives but achieve lower quality and slower speeds compared to 3DGS. Blanc et al. (2024) enables differentiable ray casting of irregularly distributed Gaussians using a BVH structure, but their rendering is slower than rasterization-based methods. Moenne-Loccoz et al. (2024) performs ray tracing with BVH for secondary lighting effects such as shadows and reflections, but this approach is approximately three times slower than rasterization. Zhou et al. (2024a) proposes a unified rendering primitive based on 3D Gaussian distributions, enabling physically based scattering for accurate global illumination but do not achieve real-time performance. In contrast, our 6DGS method slices 6D Gaussians into conditional 3D Gaussians and renders via rasterization, approximating physically based ray-traced images with high fidelity while achieving real-time performance.
@@ 4.3 Improved Control of Gaussians
To extract the necessary scale and rotation information from our 6D representation, we utilize the conditional covariance matrix $\Sigma_{\text{cond}}$. By performing Singular Value Decomposition (SVD) on $\Sigma_{\text{cond}}$, we can decompose the matrix into its principal components, revealing both the rotation and scale of the Gaussian. The decomposition is given by $\Sigma_{\text{cond}}=USV^{\top}$, where $U$ and $V$ are orthogonal matrices, and $S$ is a diagonal matrix containing the singular values. From this decomposition: the rotation matrix $R$ is derived from the left singular vectors as $R=U$, and the scale vector $s$ is obtained by taking the square root of the diagonal elements of $S$ as $s=\sqrt{\text{diag}(S)}$. To ensure that the rotation matrix $R$ forms a right-handed coordinate system, we adjust its last column based on the sign of its determinant: $R_{:,3}=R_{:,3}\cdot\text{sign}(\text{det}(R)).$
@@
6:Perform SVD as: $\Sigma_{\text{cond}}=USV^{\top}$
7:Extract the rotation matrix and scale: $R=U,\quad s=\sqrt{\operatorname{diag}(S)},\quad R_{:,3}=R_{:,3}\cdot\operatorname{sign}(\det(R))$
@@ 5.1 Experimental Protocol
We evaluate 6DGS on two datasets: the public Synthetic NeRF dataset (Mildenhall et al., 2020) and a custom dataset rendered using physically-based ray tracing (PBRT), which we refer to as the 6DGS-PBRT dataset. The 6DGS-PBRT dataset consists of six scenes: 1) cloud from the Walt Disney Animation Studios volumetric cloud dataset  [^1] https://disneyanimation.com/resources/clouds/; 2) bunny-cloud, explosion, and smoke from OpenVDB volumetric models [^2] https://www.openvdb.org/download/; 3) suzanne, the standard Blender test mesh, with a “Glasss BSDF" translucent material applied to it; 4) ct-scan, prepared from a real CT scan.
We rendered these scenes in Blender using its PBRT engine “Cycle". At the image sizes (with width equal to height) as listed in Table 2 on an NVIDIA Tesla V100 GPU, the rendering times per view were as follows: bunny-cloud took 504.0 seconds per view, cloud 838.6 seconds, explosion 35.5 seconds, smoke 71.5 seconds, suzanne 9.1 seconds, and ct-scan 28.5 seconds. For the ct-scan object, we generated 360 views with corresponding camera poses and randomly selected 324 images for training and 36 for testing. For each of the other objects, we rendered 150 views, randomly selecting 100 images for training and 50 for testing. We will make our 6DGS-PBRT dataset publicly available to the community.
@@
In our experiments, we set $\lambda_{\text{opa}}=0.35$ and the minimum opacity threshold $\tau=0.01$. For learnable $\lambda_{\text{opa}}$, we initialize $\lambda_{\text{opa}}=0.35$ and make it trainable only during the iterations of 150,000 - 280,000. All other parameters are set to their default values as in 3DGS (Kerbl et al., 2023). For the ct-scan object, we initialize the point cloud using the marching cubes algorithm as in DDGS (Gao et al., 2024). For the other objects and the Synthetic NeRF dataset (Mildenhall et al., 2020), we randomly initialize the point cloud with 100,000 points within a cube encompassing the scene. Training is performed on a single NVIDIA Tesla V100 GPU with 16 GB of memory, using the Adam optimizer (Kingma & Ba, 2014). We set the learning rate to $1\times 10^{-2}$ for the 6D covariance parameters and $1\times 10^{-3}$ for the direction component $\mu_{d}$. The default learning rates from 3DGS are applied to the remaining parameters.
Figure 3: Qualitative comparison of methods on the 6DGS-PBRT dataset (zoom in for details).
@@ 5.2 Comparison with State-of-the-Art
Table 1: Comparison of methods on the 6DGS-PBRT dataset.
@@
Table 1 compares our 6DGS method with 3DGS and N-DG on the 6DGS-PBRT dataset. Our 6DGS achieves significantly better image quality, with an average improvement of +10.08 dB in PSNR, while using significantly fewer Gaussian points (a reduction of 41.4% on average) compared to 3DGS. Compared to N-DG, our 6DGS also achieves higher image quality (+1.51 dB on average) and uses significantly fewer Gaussian points (a reduction of 82.8% on average).
Table 2: Rendering speed (FPS) comparison of methods on the 6DGS-PBRT dataset.
@@ 5.3 Ablation Study
We conduct ablation experiments on both the 6DGS-PBRT dataset and Synthetic NeRF dataset (see Appendix Table 5) to investigate the effects of various components in our 6DGS method. Specifically, we examine the impact of the parameter $\lambda_{\text{opa}}$ in the conditional probability density function (PDF) of the directional component (see Appendix Figure 4), defined as $f_{\text{cond}}=\exp(-\lambda_{\text{opa}}\cdot D)$, where $D=(d-\mu_{d})^{\top}\Sigma_{d}^{-1}(d-\mu_{d})$ represents the Mahalanobis distance between the viewing direction $d$ and the Gaussian mean direction $\mu_{d}$, and $\lambda_{\text{opa}}$ is a scalar between 0 and 1 that controls the influence of the view direction on opacity. Our ablation experiments include several settings related to $\lambda_{\text{opa}}$:
@@
Table 4: Ablation study on the 6DGS-PBRT dataset.
@@ 6 Conclusion
In this work, we introduce 6D Gaussian splatting (6DGS) that builds on the foundations laid by 3DGS and N-dimensional Gaussians (N-DG). By improving the handling of color and opacity within the 6D spatial-angular framework and optimizing the adaptive control of Gaussians using additional directional information, we develop a method that not only maintains compatibility with the 3DGS framework but also offers superior rendering capabilities, particularly in modeling complex view-dependent effects. Our extensive experiments on the custom physically-based ray tracing dataset (6DGS-PBRT) demonstrate the effectiveness of 6DGS in achieving higher image quality and faster rendering speed compared to 3DGS and N-DG.

### v3

@@ Abstract
Novel view synthesis has advanced significantly with the development of neural radiance fields (NeRF) and 3D Gaussian splatting (3DGS). However, achieving high quality without compromising real-time rendering remains challenging, particularly for physically-based rendering using ray/path tracing with view-dependent effects. Recently, N-dimensional Gaussians (N-DG) introduced a 6D spatial-angular representation to better incorporate view-dependent effects, but the Gaussian representation and control scheme are sub-optimal. In this paper, we revisit 6D Gaussians and introduce 6D Gaussian Splatting (6DGS), which enhances color and opacity representations and leverages the additional directional information in the 6D space for optimized Gaussian control. Our approach is fully compatible with the 3DGS framework and significantly improves real-time radiance field rendering by better modeling view-dependent effects and fine details. Experiments demonstrate that 6DGS significantly outperforms 3DGS and N-DG, achieving up to a 15.73 dB improvement in PSNR with a reduction of 66.5% Gaussian points compared to 3DGS. The project page is: https://gaozhongpai.github.io/6dgs/.
Figure 1: Visualizations of volumetric rendering. Top-left: our 6DGS rendering; bottom-right: physically-based rendering using ray/path tracing; right: comparison with 3DGS over the red regions.
@@ 1 Introduction
We validate our method through extensive experiments on two datasets: a custom dataset with physically based rendering using ray tracing, named the 6DGS-PBR dataset, and a public dataset without strong view-dependent effects, i.e., the Synthetic NeRF dataset (Mildenhall et al., 2020). The 6DGS-PBR dataset contains scenes with complex geometries, materials, and lighting conditions that exhibit strong view-dependent effects, making it suitable for testing our method’s capabilities in handling complex light interactions. Our results demonstrate that 6DGS significantly outperforms existing methods in both rendering quality and efficiency on the 6DGS-PBR dataset and achieves comparable performance on the Synthetic NeRF dataset. Specifically, we achieve up to a 15.73 dB improvement in peak signal-to-noise ratio (PSNR) while using only 33.5% of the Gaussian points compared to 3DGS. These improvements underscore the effectiveness of our approach in capturing fine details and complex view-dependent effects in real-time rendering scenarios.
@@
  • Validation: We validate our 6DGS method on a custom physically-based ray tracing dataset (6DGS-PBR), demonstrating its superiority in both image quality and rendering speed compared to existing approaches. Additionally, our evaluation on a public dataset showcases its generalizability to scenes without strong view-dependent effects.
@@ 2 Related Work
3D Gaussian Splatting. 3D Gaussian splatting (3DGS) (Kerbl et al., 2023) has emerged as a significant advancement in computer graphics and 3D vision, achieving high-fidelity rendering quality while maintaining real-time performance. Numerous works have been proposed to improve rendering quality (Yu et al., 2024; Lu et al., 2024), rendering efficiency (Lee et al., 2024; Bagdasarian et al., 2024), and training optimization (Kheradmand et al., 2024; Höllein et al., 2024), as well as to explore applications (Kocabas et al., 2024; Zhou et al., 2024b; Niedermayr et al., 2024) and extensions (Charatan et al., 2024; Luiten et al., 2024; Wu et al., 2024; Tang et al., 2024; Gao et al., 2024) of 3DGS. For example, Mip-Splatting (Yu et al., 2024) introduces a Gaussian low-pass filter based on Nyquist’s theorem to address aliasing and dilation artifacts by matching the maximal sampling rate across all observed samples. Compact3D (Lee et al., 2024) applies vector quantization to compress different attributes into corresponding codebooks, storing the index of each Gaussian to reduce storage overhead. 3DGS-MCMC (Kheradmand et al., 2024) proposes densification and pruning strategies in 3DGS as deterministic state transitions of Markov Chain Monte Carlo (MCMC) samples instead of using heuristics. Many of these advancements and applications of 3DGS can potentially be applied to our 6DGS, which extends 3DGS by incorporating an additional directional component.
@@
N-dimensional Gaussians. To enhance the 3D Gaussian representation, researchers have introduced other Gaussian representations for rendering. 2DGS (Huang et al., 2024) introduces a perspective-accurate 2D splatting process utilizing ray-splat intersection and rasterization to enhance geometry reconstruction. 3D-HGS (Li et al., 2024) proposes a 3D half-Gaussian kernel to improve performance without compromising rendering speed. 4DGS (Yang et al., 2024) proposes to approximate the underlying spatio-temporal 4D volume of a dynamic scene by optimizing a collection of 4D primitives. N-DG (Diolatzis et al., 2024) introduces N-dimensional Gaussians (N-DG) along with a high-dimensional culling scheme inspired by locality-sensitive hashing. Specifically, N-DG introduces a 10-dimensional Gaussian (10-DG) that incorporates geometry and material information such as world position, view direction, albedo, and roughness, as well as a 6-dimensional Gaussian (6-DG) that includes world position and view direction. Our 6DGS combines the strengths of 3DGS and N-DG for better representation and better adaptive control of Gaussians.
@@
Gaussian-based Ray Tracing. Unlike most 3DGS methods that render Gaussians via rasterization, some approaches have proposed Gaussian-based ray tracing, generally at the cost of slower rendering speeds or even lower quality. For instance, Condor et al. (2024) models scattering and emissive media using mixtures of simple kernel-based volumetric primitives but achieve lower quality and slower speeds compared to 3DGS. Blanc et al. (2024) enables differentiable ray casting of irregularly distributed Gaussians using a BVH structure, but their rendering is slower than rasterization-based methods. Moenne-Loccoz et al. (2024) performs ray tracing with BVH for secondary lighting effects such as shadows and reflections, but this approach is approximately three times slower than rasterization. Zhou et al. (2024a) proposes a unified rendering primitive based on 3D Gaussian distributions, enabling physically based scattering for accurate global illumination but do not achieve real-time performance. In contrast, our 6DGS method slices 6D Gaussians into conditional 3D Gaussians and renders via rasterization, approximating physically based ray-traced images with high fidelity while achieving real-time performance.
@@ 4.3 Improved Control of Gaussians
To extract the necessary scale and rotation information from our 6D representation, we utilize the conditional covariance matrix $\Sigma_{\text{cond}}$. By performing Singular Value Decomposition (SVD) on $\Sigma_{\text{cond}}$, we can decompose the matrix into its principal components, revealing both the rotation and scale of the Gaussian. The decomposition is given by $\Sigma_{\text{cond}}=UDU^{\top}$, where $U$ is an orthogonal matrix, and $D$ is a diagonal matrix containing the singular values. From this decomposition: the rotation matrix $R$ is derived from the left singular vectors as $R=U$, and the scale vector $S$ is obtained by taking the square root of the diagonal elements of $D$ as $S=\sqrt{\text{diag}(D)}$. To ensure that the rotation matrix $R$ forms a right-handed coordinate system, we adjust its last column based on the sign of its determinant: $R_{:,3}=R_{:,3}\cdot\text{sign}(\text{det}(R)).$
@@
6:Perform SVD as: $\Sigma_{\text{cond}}=UDU^{\top}$
7:Extract rotation matrix and scale: $R=U,\quad S=\sqrt{\operatorname{diag}(D)},\quad R_{:,3}=R_{:,3}\cdot\operatorname{sign}(\det(R))$
@@ 5.1 Experimental Protocol
We evaluate 6DGS on two datasets: the public Synthetic NeRF dataset (Mildenhall et al., 2020) and a custom dataset using physically-based rendering (PBR), which we refer to as the 6DGS-PBR dataset. The 6DGS-PBR dataset consists of six scenes: 1) cloud from the Walt Disney Animation Studios volumetric cloud dataset  [^1] https://disneyanimation.com/resources/clouds/; 2) bunny-cloud, explosion, and smoke from OpenVDB volumetric models [^2] https://www.openvdb.org/download/; 3) suzanne, the standard Blender test mesh, with a “Glasss BSDF” translucent material applied to it; 4) ct-scan, prepared from a real CT scan.
We rendered these scenes in Blender using its PBR engine “Cycle”. At the image sizes (with width equal to height) as listed in Table 2 on an NVIDIA Tesla V100 GPU, the rendering times per view were as follows: bunny-cloud took 504.0 seconds per view, cloud 838.6 seconds, explosion 35.5 seconds, smoke 71.5 seconds, suzanne 9.1 seconds, and ct-scan 28.5 seconds. For the ct-scan object, we generated 360 views with corresponding camera poses and randomly selected 324 images for training and 36 for testing. For each of the other objects, we rendered 150 views, randomly selecting 100 images for training and 50 for testing. We will make our 6DGS-PBR dataset publicly available to the community.
@@
In our experiments, we set $\lambda_{\text{opa}}=0.35$ and the minimum opacity threshold $\tau=0.01$. For learnable $\lambda_{\text{opa}}$, we initialize $\lambda_{\text{opa}}=0.35$ and make it trainable only during the iterations of 15,000 - 28,000. All other parameters are set to their default values as in 3DGS (Kerbl et al., 2023). For the ct-scan object, we initialize the point cloud using the marching cubes algorithm as in DDGS (Gao et al., 2024). For the other objects and the Synthetic NeRF dataset (Mildenhall et al., 2020), we randomly initialize the point cloud with 100,000 points within a cube encompassing the scene. Training is performed on a single NVIDIA Tesla V100 GPU with 16 GB of memory, using the Adam optimizer (Kingma & Ba, 2014). We set the learning rate to $1\times 10^{-2}$ for the 6D covariance parameters and $1\times 10^{-3}$ for the direction component $\mu_{d}$. The default learning rates from 3DGS are applied to the remaining parameters.
Figure 3: Qualitative comparison of methods on the 6DGS-PBR dataset (zoom in for details).
@@ 5.2 Comparison with State-of-the-Art
Table 1: Comparison of methods on the 6DGS-PBR dataset.
@@
Table 1 compares our 6DGS method with 3DGS and N-DG on the 6DGS-PBR dataset. Our 6DGS achieves significantly better image quality, with an average improvement of +10.08 dB in PSNR, while using significantly fewer Gaussian points (a reduction of 41.4% on average) compared to 3DGS. Compared to N-DG, our 6DGS also achieves higher image quality (+1.51 dB on average) and uses significantly fewer Gaussian points (a reduction of 82.8% on average).
Table 2: Rendering speed (FPS) comparison of methods on the 6DGS-PBR dataset.
@@ 5.3 Ablation Study
We conduct ablation experiments on both the 6DGS-PBR dataset and Synthetic NeRF dataset (see Appendix Table 8) to investigate the effects of various components in our 6DGS method. Specifically, we examine the impact of the parameter $\lambda_{\text{opa}}$ in the conditional probability density function (PDF) of the directional component (see Appendix Figure 5), defined as $f_{\text{cond}}=\exp(-\lambda_{\text{opa}}\cdot D)$, where $D=(d-\mu_{d})^{\top}\Sigma_{d}^{-1}(d-\mu_{d})$ represents the Mahalanobis distance between the viewing direction $d$ and the Gaussian mean direction $\mu_{d}$, and $\lambda_{\text{opa}}$ is a scalar between 0 and 1 that controls the influence of the view direction on opacity. Our ablation experiments include several settings related to $\lambda_{\text{opa}}$:
@@
Table 4: Ablation study on the 6DGS-PBR dataset.
@@ 6 Conclusion
In this work, we introduce 6D Gaussian splatting (6DGS) that builds on the foundations laid by 3DGS and N-dimensional Gaussians (N-DG). By improving the handling of color and opacity within the 6D spatial-angular framework and optimizing the adaptive control of Gaussians using additional directional information, we develop a method that not only maintains compatibility with the 3DGS framework but also offers superior rendering capabilities, particularly in modeling complex view-dependent effects. Our extensive experiments on the custom physically-based rendering dataset (6DGS-PBR) demonstrate the effectiveness of 6DGS in achieving higher image quality and faster rendering speed compared to 3DGS and N-DG.
@@ Appendix A Appendix
A.1 Real-world Scenes
Table 5: Comparison of methods on real-world datasets: Deep Blending (Hedman et al., 2021) and Tanks & Temples (Knapitsch et al., 2017).
Dataset Scene 3DGS N-DG 6DGS (Ours)
PSNR SSIM # points PSNR SSIM # points PSNR SSIM # points
Deep Blending
drjohnson 29.22 0.898 3,276,989 26.31 0.828 494,494 28.12 0.883 2,074,490
playroom 29.74 0.900 2,332,830 27.74 0.866 258,999 29.06 0.892 1,685,626
avg 29.48 0.899 2,804,910 27.03 0.847 376,747 28.59 0.888 1,880,058
Tanks& Temples train 21.75 0.803 1,100,525 13.26 0.440 996,826 21.95 0.787 839,408
truck 25.08 0.869 2,606,855 13.65 0.462 663,616 25.05 0.859 2,050,162
avg 23.42 0.836 1,853,690 13.46 0.451 830,221 23.50 0.823 1,444,785
Table 6: Comparison of methods on the real-world dataset: Shiny (Wizadwongsa et al., 2021)
. Scene 3DGS 6DGS (Ours) PSNR SSIM # points PSNR SSIM # points cd 25.51 0.843 1,128,098 28.39 0.895 596,789 crest 18.80 0.622 4,611,724 19.35 0.648 3,144,178 food 18.39 0.500 2,362,888 17.97 0.475 1,190,056 giants 24.24 0.844 2,341,337 24.15 0.826 1,635,315 lab 24.69 0.836 843,202 27.66 0.903 490,878 pasta 15.45 0.373 2,287,582 14.90 0.349 978,188 seasoning 26.25 0.823 1,085,732 26.36 0.811 524,848 tools 26.20 0.908 1,180,973 25.28 0.884 593,190 avg 22.44 0.719 1,980,192 23.01 0.724 1,144,180
We evaluate 6DGS on three real-world datasets: Deep Blending (Hedman et al., 2021), Tanks & Temples (Knapitsch et al., 2017), and Shiny (Wizadwongsa et al., 2021) to validate its robustness and effectiveness in practical scenarios, as shown in Table 5 and 6. These datasets present diverse challenges, including scenes with varying levels of complexity, strong view-dependent effects, and detailed geometric structures.
In the Deep Blending dataset (Hedman et al., 2021), 6DGS achieves comparable quality to 3DGS, with an average PSNR of 28.59 versus 29.48, while reducing the number of Gaussian points by 33%. This demonstrates that 6DGS effectively maintains high-quality rendering even with significantly fewer points, thereby improving computational efficiency without largely compromising accuracy.
On the Tanks & Temples dataset (Knapitsch et al., 2017), which features large-scale outdoor scenes with intricate geometric details, 6DGS outperforms 3DGS with a slight improvement in PSNR (23.50 vs. 23.42) and a 22% reduction in the number of points used. These results highlight the ability of 6DGS to generalize effectively to outdoor scenes and achieve improved efficiency.
The Shiny dataset (Wizadwongsa et al., 2021), known for its challenging scenes with strong view-dependent effects, offers a rigorous test for 6DGS’s directional modeling capabilities. Following the settings in NeX (Wizadwongsa et al., 2021), we resize the images to $1008\times 567$ for cd and lab and to $1008\times 756$ for other scenes. In this dataset, 6DGS excels, particularly in scenes such as cd and lab, where it achieves PSNR gains of +2.88 and +2.97, respectively, compared to 3DGS. Additionally, these gains are achieved while utilizing approximately 50% fewer Gaussian points. These results underscore the strength of 6DGS in handling view-dependent phenomena.
Across all datasets, the improvements in PSNR and point efficiency demonstrate the robustness and versatility of 6DGS in real-world settings. By reducing the number of Gaussian points while maintaining or exceeding rendering quality, 6DGS offers a significant advancement in computational efficiency and scalability for practical applications. This ability to model strong view-dependent effects and detailed geometry efficiently positions 6DGS as a valuable tool for various rendering tasks, from real-time applications to high-fidelity visualizations.
A.2 Subsurface Scattering Scene
Figure 4: Qualitative comparison of methods on the subsurface scatter (SSS) dragon scene.
Table 7: Comparison of methods on the scene of subsurface scattering (SSS) dragon.
SSS Dragon PSNR SSIM # points FPS Train (min)
3DGS 26.57 0.813 269,250 104.6 24
N-DG 33.19 0.936 196,645 86.0 50
6DGS (Ours) 35.00 0.937 128,748 111.5 33
6DGS-Flash (Ours) - - - 324.3 -
To further validate the versatility of our method, we have added a new dragon scene that exhibits both subsurface scattering (SSS) and surface scattering effects. Following the same evaluation protocol as other scenes (except for ct-scan), we rendered 150 views and randomly selected 100 images for training and 50 for testing. As shown in Table 7, 6DGS significantly outperforms both baselines, achieving the best PSNR (35.00) and SSIM (0.937) while using substantially fewer Gaussian points. Our method also maintains competitive training efficiency (33 minutes vs. 24 minutes for 3DGS and 50 minutes for N-DG) while achieving real-time rendering performance (111.5 FPS, further accelerated to 324.3 FPS with 6DGS-Flash). Figure 4 demonstrates that 6DGS achieves superior visual quality and more faithful geometry reconstruction, further validating our method’s effectiveness in handling diverse scenarios involving both volumetric and surface-based light transport phenomena.
A.3 View-dependent Opacity